OpenStudy (anonymous):

MEDALS ...A flat rectangular piece of aluminum has a perimeter of 70 inches. The length is 11 inches longer than the width. Find the width.

OpenStudy (ivettef365):

formula is P = 2L + 2W

OpenStudy (anonymous):

24 thank u!

OpenStudy (anonymous):

lol well thats not a answer guess 23

OpenStudy (anonymous):

not a selection

OpenStudy (anonymous):

so weird ... its multiple choice

OpenStudy (ivettef365):

Oh, ok, it saids that L = w + 11

OpenStudy (ivettef365):

ok, so it should be 70 = 2(w+11) + 2w

OpenStudy (ivettef365):

now you can solve for w, my mistake thought L equals 11

OpenStudy (anonymous):

so 23 u got also?

OpenStudy (ivettef365):

70 = 2(w+11) + 2w 70 = 2w + 22 + 2w 70 -22 = 4w 48 = 4w w = 48/4 w = 12

OpenStudy (ivettef365):

so L = 12 + 11 = 23 W = 12 23 + 23 + 12 + 12 = 70 true

OpenStudy (anonymous):

ty for ur time

OpenStudy (ivettef365):

yw :)